# An Overview of Some Single Machine Scheduling Problems: Polynomial Algorithms, Complexity and Approximability

### v1

@@ 3.2 Minimizing the Number of Late Jobs
Next, we briefly discuss some directions for further research. In the multi-processor case, the weighted version is known to be NP-complete [12], but the non-weighted version remains open. More specifically, it is not known whether the problem $P|r_{j};p_{j}=p;{pmtn}|\sum U_{j}$ can be solved in polynomial time. (One difficulty that arises even for $2$ processors, is that we cannot restrict ourselves to earliest deadline schedules. For example, an instance consisting of three jobs with feasible intervals $(0,3)$, $(0,4)$, and $(0,5)$ and the processing time $p=3$ is feasible, but the earliest-deadline schedule will complete only jobs $1$ and $2$.) For the multi-processor case, one can also consider a preemptive version, where jobs are not allowed to migrate between processors.
@@ 3.5 Scheduling with Tardiness-based Objectives
In [35], a new general graphical algorithm was presented which often improves the complexity or at least the running time of dynamic programming algorithms, and it can be used for efficient approximation algorithm. Compared to standard dynamic programming algorithms, the number of states can often substantially be reduced.
@@ 3.6 Preemptive Scheduling in Overloaded Systems
In the single machine scheduling problem in overloaded systems, each job has a release time, a deadline and a weight. The latter parameter reflects its profit. Any job can be preempted during the execution, and the objective is to maximize the overall profit. The profit gained from processing task $j$ can be defined in two ways. In the standard model, each completed task $j$ gives the profit $w_{j}p_{j}$, but non-completed tasks do not yield any profit. In the metered model, a task $j$ executed for time $t\leq p_{j}$ gives the profit $w_{jt}$ even if it is not completed. For the off-line metered profit model, Chrobak et al. [16] used a reduction to bipartite matchings and maximal flows to develop a polynomial-time algorithm. For the online metered case, they presented an algorithm with competitive ratio $e/(e-1)\approx 1.5820$ and proved a lower lower bound of $\sqrt{5}-1\approx 1.236$ on the competitive ratio of algorithms for this problem.
@@ 3.7 Scheduling with Restarts
in which a waiting strategy is incorporated. Van den Akker, Hoogeveen and Vakhania [3] considered an extension of the standard model by allowing job restarts: if an urgent job comes in while the machine is busy, then the currently running job can be interrupted in favor of the former urgent job. In this model, unlike traditional preemptive models, the already processed part of the interrupted job is wasted, this job is to be started all over again. The algorithm proposed in [3] has the best possible worst-case ratio of $1.5$.
Again, we give some possible directions for further research in connection with allowed restarts. An interesting question is to determine the effect of a partial restart, in which a fraction of the work spent on a job whose processing is broken off is subtracted from its processing time. It would be particularly interesting to find out the relationship between the worst-case ratio and this ‘reduction percentage’ $p$; it might be a straight line between 1.5 (complete restart) and 1 (preemption allowed).
Non-strict lower bounds, which were obtained in almost linear time, can be used in any search-tree based approximation algorithm. An example of such an algorithm is beam search. Similarly, using our strict lower bounds with the solution tree $T$ from this section, we obtain exact branch-and-bound algorithms. Once implemented and tested, the above algorithms can be used as an algorithmic engine for the construction of a decision support system for flexible job shop scheduling problems. The presented algorithms can be easily aggregated by an additional graph-completion mechanism for taking into the account the transportation times and other real-life job shop scheduling problems.

### v2

@@ 3.2 Minimizing the Number of Late Jobs
Next, we briefly discuss some directions for further research. First we note that for the multi-processor case, the weighted version is $NP$-complete [12]. At the same time, the non-weighted version still remains open. In particular, it is not known whether the problem $P|r_{j};p_{j}=p;{pmtn}|\sum U_{j}$ can be solved in polynomial time. (Even for $2$ processors, one cannot the search to solely earliest deadline schedules. As an example, consider an instance with three jobs with feasible intervals $(0,3)$, $(0,4)$, and $(0,5)$ and the processing time $p=3$. There is a feasible schedule that completes all three jobs, whereas the earliest-deadline schedule completes only jobs $1$ and $2$.) For the multi-processor case, it is also interesting to study the preemptive version, where jobs are not allowed to migrate between processors.
@@ 3.5 Scheduling with Tardiness-based Objectives
In [36], Lazarev and Werner presented a new general graphical algorithm which often improves the complexity or at least the running time of dynamic programming algorithms, and it can be used for efficient approximation algorithm. Compared to standard dynamic programming algorithms, the number of states can often substantially be reduced.
@@ 3.6 Preemptive Scheduling in Overloaded Systems
In the single machine scheduling problem in overloaded systems, each job has a release time, a deadline and a weight. The latter parameter reflects its profit. Any job can be preempted during the execution, and the objective is to maximize the overall profit. Two possibly models and the definition of the profit gained by the execution of task $j$ can be defined. In the standard model, a completed task $j$ gives the profit $w_{j}p_{j}$, whereas a non-completed task gives no profit. In the metered model, a task $j$ executed for $t\leq p_{j}$ time units gives the profit $w_{jt}$ (this task may not be completed). For the off-line metered profit model, Chrobak et al. [16] used a reduction to bipartite matchings and maximal flows to develop a polynomial-time algorithm. For the online metered case, they presented an algorithm with competitive ratio $e/(e-1)\approx 1.5820$ and proved a lower lower bound of $\sqrt{5}-1\approx 1.236$ on the competitive ratio of algorithms for this problem.
@@ 3.7 Scheduling with Restarts
in which a waiting strategy is incorporated. Van den Akker, Hoogeveen and Vakhania [3] considered an extension of the standard model by allowing job restarts: if an urgent job comes in while the machine is busy, then the currently running job can be interrupted in favor of the former urgent job. In this model, unlike traditional preemptive models, the already processed part of the interrupted job is wasted, this job is to be started all over again. The algorithm proposed in [3] has the best possible worst-case ratio of $1.5$ (which is better than the above mentioned bound for the case without the restarts).
Again, we give some possible directions for further research in connection with allowed restarts. As we just mentioned, with the restart model considered in [3], if a job is restarted, then the portion of this job that was already processed by the machine is lost. This may be avoided for some applications, where the already completed part of an abandoned job gives some benefit and this job can be restarted from the time moment when it was interrupted or at some earlier time. Hence, one may consider models with partial restarts when the already processed part of an interrupted job counts partly or completely. In particular, it is interesting to find out whether there exist better worst-case performance bounds for models with partial restarts.
@@ 4 Concluding Remarks
For single machine scheduling problems, efficient solution methods can often be found easier than for the corresponding multiprocessor or shop scheduling problems. However, such methods may give a better insight into the latter more complex problems and may well serve as a basis for the development of solution methods for these problems. At the same time, optimal solutions to the single machine scheduling problems can directly be used to solve the corresponding multiprocessor or shop scheduling problems. For example, for such $NP$-hard problems, strong lower bounds can be obtained based on the solutions of the corresponding single machine problems. Non-strict lower bounds, which are easier and faster to obtain, can be used in any search-tree based approximation algorithm, such as beam search, for example. Similarly, using strict lower bounds, we obtain exact branch-and-bound algorithms. Once implemented and tested, the above algorithms can be used as an algorithmic engine for the construction of decision support systems. Finally, we note that a number of algorithms for basic single machine problems can relatively easily be adopted using additional graph-completion mechanisms to solve the corresponding extended versions with precedence relations, transportation and setup times and other real-life job scheduling problems.